Consider

f(t) = cos(9t),  g(t) = et.

Proceed as in this example and find the convolution f ∗ g of the given functions.
Answer #1

g(t)=et Here f(t) = cos(92) By definition of convolution t-) (f+9) (t) = cos(gi). e. .de t Let I t s cos(at).edt ( cos(at) et